# RL-100: Performant Robotic Manipulation with Real-World Reinforcement Learning

### v1

@@ 5.2 Description of Tasks
Following the overview, we describe the objective, challenges and further evaluation protocols beyond normal ones of each task individually.
@@
Dynamic Unscrewing (1024, 3) 23 Same as above 23 = 22 + 1 Same as above, 1: enable arm motion
@@ 5.2.2 Agile Bowling
To evaluate the generalization capabilities of the framework, we test the task under varying initial positions of the curling stone and bowling pins along their respective lines. We further test (1) zero-shot transferability on a coarser trail surface and (2) adaptation by further finetuning on an inversed pins placement of the policy. These variations challenge the policy to adapt its pushing strategy and trajectory corrections to different configurations while maintaining accuracy.
@@ 5.2.6 Orange juicing
The orange juicing task consists of three subtasks: placing a halved orange on the juicer, pressing down the lever to extract juice, and removing the discarded orange pulp. The robot’s end effector is replaced with two 3D-printed sleeves that fix two rubber grippers, enabling robust grasping irregular fruit surfaces. For each subtask, the initial and final poses of the robot arm are fixed. Note that these poses are defined individually for each subtask and therefore differ across the three subtasks.
[…]
Lever pressing. The pressing process follows a relatively fixed trajectory. To ensure consistent execution, we use kinesthetic teaching to record the downward and upward motions of the lever in advance. During deployment, the robot replays this trajectory to lower the hammer-like press, squeeze the juice out of the orange, and then reset the lever. The task is considered successful if the lever is fully pressed down and restored.
Table 3: Success rate (%) across tasks. RL–100 (ours) groups an iterative offline fine-tuning stage followed by online RL with either DDIM or one-step CM policy. Averages are unweighted across tasks.
@@
 Mean over the six tasks evaluated with CM.
@@ 6.2 Zero-shot adaptation to new dynamics
To assess the robustness of our learned policies, we evaluate zero-shot transfer to environmental variations unseen during training, listed in Tab. 4 and Fig. 8. On Dynamic Push-T, as shown in Fig. 8B-C, the policy maintains perfect 100% success when the surface friction is substantially altered, and achieves 80% success even when visual and physical clutter from interference objects of various shapes are introduced into the workspace. Similarly, Agile Bowling in Fig. 8D achieves 100% success on modified surface properties that change ball rolling dynamics. For Pouring in Fig. 8A, replacing granular nuts with fluid water—a significant shift in material properties and flow dynamics—still yields 90% success, demonstrating that the policy has learned sufficiently stable and robust manipulation strategies rather than overfitting to specific physical parameters. Across all four variations, our method achieves 92.5% average success without any retraining or fine-tuning, indicating strong generalization to distribution shifts in environmental dynamics.
@@ 7 Simulation Experiments
To complement our real-robot evaluations, we benchmark RL-100 in simulation and compare it against state-of-the-art diffusion/flow–based offline-to-online RL methods, including DPPO (Ren et al., 2024), ReinFlow (Zhang et al., 2025), and DSRL (Wagenmaker et al., 2025). We also conduct ablations to verify the contribution of each component in RL-100.
@@ 7.2 Ablation Studies
We conduct ablation studies on five components: (i) visual modality (2D vs. 3D), (ii) diffusion-noise clipping, (iii) sampler/model (CM vs. DDIM), (iv) representation learning during fine-tuning (ReconVIB / no ReconVIB / frozen encoder), and (v) diffusion parameterization ($\epsilon$- vs. $x_{0}$-prediction). The comparison is listed in Fig. 15 and the detailed configurations and results are provided below. Visual modality: 2D vs. 3D. Our framework is modality-agnostic and accepts either 2D RGB images or 3D point clouds as visual inputs. On the Adroit door task—a relatively clean scene—the 3D variant learns faster and attains a higher final success rate. We attribute this to the ability of 3D inputs to enable precise geometric filtering (e.g., ROI/radius “clipping”) that cleanly isolates task-relevant structures such as the handle and contact surfaces, yielding a higher signal-to-noise ratio and easier credit assignment than 2D images. This ablation thus demonstrates both the flexibility of our framework (2D/3D optional) and the practical advantage of 3D in clean, contact-rich settings.
@@ 8 Conclusion and Future Work
Beyond excelling under nominal conditions, RL-100 exhibited strong generalization and robustness in the face of new variations and disturbances. Without any retraining, a single policy maintained high performance under significant environmental shifts. Across several such unseen scenario variations, RL-100 attained an average 92.5% success rate without any fine-tuning, underscoring the policy’s robustness to modest domain shifts. Furthermore, the framework’s consistency-model variant (the one-step distilled policy) demonstrated that high-frequency control is possible without loss of accuracy, as it too reached a 100% success rate on all evaluated tasks, including 250/250 successes in the challenging dual-arm towel folding. This blend of reliability and adaptability indicates that RL-100 can handle the kind of dynamic perturbations and variability inherent in real home or factory environments, rather than overfitting to narrow training conditions.

### v2

@@ 5.2 Description of Tasks
Following the overview, we describe the objective, challenges, and further evaluation protocols beyond the normal ones of each task individually.
@@
Dynamic Unscrewing (1024, 3) 23 Same as above 23 = 22 + 1 22: Same as above, 1: enable arm motion
@@ 5.2.2 Agile Bowling
To evaluate the generalization capabilities of the framework, we test the task under varying initial positions of the curling stone and bowling pins along their respective lines. We further test (1) zero-shot transferability on a coarser trial surface and (2) adaptation by further finetuning on an inversed pins placement of the policy. These variations challenge the policy to adapt its pushing strategy and trajectory corrections to different configurations while maintaining accuracy.
@@ 5.2.6 Orange juicing
The orange-juicing task comprises three subtasks: placing a halved orange on the juicer, pressing the lever to extract juice, and removing the discarded pulp. The robot’s end effector is fitted with two 3D-printed sleeves that secure rubber grippers, enabling robust grasps on irregular fruit surfaces. For each subtask, the robot arm’s initial and final poses are fixed; these poses are defined per subtask and therefore differ across the three subtasks.
Orange placing. A randomly sized half-orange is placed at a random position and orientation within a  $29\,\mathrm{cm}\times 20\,\mathrm{cm}$ inner area of a tray. The robot first grasps the orange with the gripper oriented vertically to the tray surface, then rotates the wrist joint by $180^{\circ}$ to flip the orange, and finally places it onto the filter section of a lever-style juicer before retracting. Success is defined as the orange being stably seated on top of the filter. This step is challenging due to large variability in size, inclination, height, and placement (see Fig. 7A and Fig. 7B), requiring strong robustness to spatial randomness.
Lever pressing. The pressing motion follows a fixed trajectory. To ensure consistent execution, we record the downward and upward lever motions via kinesthetic teaching. At deployment, the robot replays this trajectory to lower the press, squeeze the juice, and reset the lever. Success is defined as fully depressing and restoring the lever.
Table 3: Success rate (%) across tasks. RL-100 (ours) groups an iterative offline fine-tuning stage followed by online RL with either DDIM or one-step CM policy. Averages are unweighted across tasks.
@@
 Mean over the six tasks evaluated with CM. Juicing - Removal exhibits IK-induced pose discontinuities between push and grasp in a tight, slippery contact; this issue is already present in the demonstrations and is exacerbated by the noise-sensitive one-step CM policy; for safety, we did not evaluate CM on this task.
@@ 6.2 Zero-shot adaptation to new dynamics
To assess the robustness of our learned policies, we evaluate zero-shot transfer to environmental variations unseen during training, listed in Tab. 4 and Fig. 8. On Dynamic Push-T, as shown in Fig. 8B-C, the policy maintains perfect 100% success when the surface friction is substantially altered, and achieves 80% success even when visual and physical clutter from interference objects of various shapes are introduced into the workspace. Similarly, Agile Bowling in Fig. 8D achieves 100% success on modified surface properties that change ball rolling dynamics. For Pouring in Fig. 8A, replacing granular nuts with fluid water—a significant shift in material properties and flow dynamics—still yields 90% success, demonstrating that the policy has learned sufficiently stable and robust manipulation strategies rather than overfitting to specific physical parameters. For soft-towel folding on unseen shape in Fig. 8E, RL-100 attains 80% success, suggesting insensitivity to towel geometry and robust representation learning. Across all four variations, our method achieves 90.0% average success without any retraining or fine-tuning, indicating strong generalization to distribution shifts in environmental dynamics.
@@ 6.6 Training efficiency
As shown in Tab. LABEL:tab:training_efficiency, our pipeline reaches deployment-grade performance with minimal human teleoperation and moderate policy rollouts. Across seven tasks, we average only 115 human demos ($\sim$1.8 h) per task, after which iterative offline RL adds $\sim$566 rollouts (6.5 h) and online RL adds $\sim$434 rollouts (5.6 h) per task. Summed over all tasks, human collection is just 804 episodes in 12.5 h—about 13% of the total << 100 h data-collection budget—while the remaining time comes from autonomous policy rollouts (offline: 3965 eps/45.3 h; online: 3037 eps/39.5 h). This balance demonstrates that a small amount of expert data, complemented by moderate, inexpensive rollouts, suffices to train policies that meet our deployment criteria.
@@
6.7 Robots Against and For Humans
Figure 13: RL-100 in competitive and service settings. (A) Bowling matches between RL-100 and several human players. (B) Public deployment in a shopping mall where RL-100 prepares fresh juice for visitors during a multi-hour demo. (C) Head-to-head comparison on a T-shaped block-pushing benchmark; RL-100 achieves more trials within the same time window.
Fig. 13 positions RL-100 relative to humans in both competitive and assistive settings, consistent with our deployment criteria (reliability, efficiency, robustness). In tabletop bowling (Fig. 13A), with 25 attempts per side, RL-100 achieves 25 successes versus 14 for human players. In a public shopping-mall demonstration (Fig. 13B), RL-100 prepares fresh juice continuously for seven hours without failure, evidencing robustness outside the lab and zero-shot transfer to a new environment. On the standardized T-push benchmark (Fig. 13C), under an equal wall-clock budget the robot completes 20 successful trials compared with 17 by an expert teleoperator and 13 by several novices, indicating lower time-to-completion. Together, these results illustrate that starting from strong human priors and post-training with real-world RL enables RL-100 to match-and in several tasks exceed-skilled human performance while remaining suitable for real deployments.
@@ 7 Simulation Experiments
To complement our real-robot evaluations, we benchmark RL-100 in simulation and compare it against state-of-the-art diffusion/flow–based offline-to-online RL methods, including DPPO (Ren et al., 2024), ReinFlow (Zhang et al., 2025b), and DSRL (Wagenmaker et al., 2025). We also conduct ablations to verify the contribution of each component in RL-100.
@@ 7.2 Ablation Studies
We conduct ablation studies on five components: (i) visual modality (2D vs. 3D), (ii) diffusion-noise clipping, (iii) sampler/model (CM vs. DDIM), (iv) representation learning during fine-tuning (ReconVIB / no ReconVIB / frozen encoder), and (v) diffusion parameterization ($\epsilon$- vs. $x_{0}$-prediction). The comparison is listed in Fig. 16 and the detailed configurations and results are provided below. Visual modality: 2D vs. 3D. Our framework is modality-agnostic and accepts either 2D RGB images or 3D point clouds as visual inputs. On the Adroit door task—a relatively clean scene—the 3D variant learns faster and attains a higher final success rate. We attribute this to the ability of 3D inputs to enable precise geometric filtering (e.g., ROI/radius “crop”) that cleanly isolates task-relevant structures such as the handle and contact surfaces, yielding a higher signal-to-noise ratio and easier credit assignment than 2D images. This ablation thus demonstrates both the flexibility of our framework (2D/3D optional) and the practical advantage of 3D in clean, contact-rich settings.
@@ 8 Conclusion and Future Work
Beyond excelling under nominal conditions, RL-100 exhibited strong generalization and robustness in the face of new variations and disturbances. Without any retraining, a single policy maintained high performance under significant environmental shifts. Across several such unseen scenario variations, RL-100 attained an average 90.0% success rate without any fine-tuning, underscoring the policy’s robustness to modest domain shifts. Furthermore, the framework’s consistency-model variant (the one-step distilled policy) demonstrated that high-frequency control is possible without loss of accuracy, as it too reached a 100% success rate on all evaluated tasks, including 250/250 successes in the challenging dual-arm towel folding. This blend of reliability and adaptability indicates that RL-100 can handle the kind of dynamic perturbations and variability inherent in real home or factory environments, rather than overfitting to narrow training conditions.